# Rethinking Self-training for Semi-supervised Landmark Detection: A Selection-free Approach

### v1

@@ V-C1 Semi-supervised Setting
In addition to the existing semi-supervised landmark detection models [50, 21, 53, 55, 56], we also compare with semi-supervised pose estimation methods [32, 51] and generic self-training methods [15, 19, 28, 66]. For selection-based methods, we select their optimal thresholds based on the performance on 300W with 5% labeled data, which can be found in Fig. 8. Following [56], we also equip our model with HRNet [47] for better performance. Tab. I shows the results on 300W, WFLW, and AFLW with different labeled ratios. We can see that STLD obtains the best results on most of the entries. Specifically, STLD-HM-R18 and STLD-TF-R18 both outperform their ResNet-18 counterparts on almost all the entries, demonstrating the effectiveness of the proposed method. When a heavier backbone such as HRNet is equipped, the performance of STLD is further boosted. For example, STLD-HM-HRNet achieves 3.73, 3.40, and 3.32 of NME (%) on 300W for 1.6%, 5%, and 10% labeled data, respectively, outperforming all the previous methods. On WFLW and AFLW, STLD-HM-HRNet achieves the best results with 5% labeled data. In contrast, STLD-TF-HRNet performs better with higher labeled ratios, thanks to the representation capability of transformers. For example, it obtains 3.25 of NME (%) on 300W and 1.53 on AFLW with 20% labeled, both of which are slightly better than STLD-HM-HRNet. The advantage of STLD-TF-HRNet is even larger on WFLW, achieving 4.46 with 20% labeled, which is about 5.7% improvement over the best existing method HybridMatch (4.73). To make the evaluation comprehensive, we also include the results under AUC and FR, which are shown in Tab. II. Again, our STLD with HRNet obtains the SOTA performance on both metrics, validating the superiority of the proposed method.
| Method | Backbone |
Unlabeled |
MRE (mm) |
| SCN [71] | UNet |
0.66 |
| GU2Net [4] | UNet |
0.84 |
| HM [64] | ResNet-50 |
0.68 |
| TF [24] | ResNet-50 |
0.69 |
| Ours w/ HM | ResNet-50 |
12K |
0.64 |
| Ours w/ TF | ResNet-50 |
12K |
0.61 |
@@ V-D Model Analysis
| Method |
1.6% |
5% |
10% |
20% |
| No Pretrain |
4.26 |
3.64 |
3.48 |
3.34 |
| MAE [76] |
4.19(-1.6%) |
3.60(-1.1%) |
3.45(-0.9%) |
3.33(-0.3%) |
| PP (Ours) |
3.97(-6.8%) |
3.47(-4.7%) |
3.39(-2.6%) |
3.30(-1.2%) |
@@
How pseudo pretraining works. One may worry that the noisy samples learned during pseudo pretraining would be memorized by the model and thus leads to confirmation bias. To understand how it works, we perform an analysis of example forgetting by computing the difference of the pseudo-labels used for training and their predictions after the two-stage learning. Note that shrink regression is not used in the second stage here as we would like to focus on pseudo pretraining. As a comparison, we also do the analysis for the naive baseline. Fig. 10 shows the relevant results of both HM (top) and TF (bottom), with different labeled ratios, and the pseudo-labels are grouped by noise. First of all, we can see that the naive baseline forgets more about noisier samples, which is consistent to the findings from [77]. And as expected, the model with the two-stage training does not memorize the pseudo-labels as much as the naive baseline does because the pseudo-labels are not used in the second stage of the proposed method. But interestingly, the forgetting of noisier samples is enlarged when compared to the naive baseline (i.e., the pink area222The green dotted line of the pink area is obtained by moving the upper green solid line down and overlaps with the lower green solid line at the first black dot), indicating that the two-stage design with pseudo pretraining is able to handle noisy pseudo-labels implicitly. Such a characteristic helps pseudo pretraining to effectively utilize the task-specific knowledge of pseudo-labels while avoiding the influence of noisy samples. To show the superiority of pseudo pretraining, we compare it with masked autoencoder (MAE) [76], a popular self-supervised pretraining method. For reference, we add a baseline named No Pretrain, which is directly fine-tuned on labeled data. Tab. V gives the results. We can see that pseudo-pretraining outperforms MAE consistently across all the labeled ratios, verifying the advantage of utilizing pseudo-labels over unlabeled pretraining as the former incorporates task-specific knowledge for better performance.
@@
How shrink loss works. While the shrink regression for heatmap-based models looks straightforward, the shrink loss applied to coordinate-based models seems not so obvious on adjusting task confidence. To verify it, we perform a correlation analysis of the gradients between GTs and pseudo-labels. We record the gradients333The seeds are fixed to make sure each pair of gradients from GTs and pseudo-labels is generated by the same batch of training data. of the last layer when trained with GTs and pseudo-labels, respectively, and group them by pseudo-label noise. For each group, we compute the Pearson correlation of the gradients between GTs and pseudo-labels. Fig. 11a and 11b show the results on 5% and 10% labeled settings of 300W, respectively. We can see that the correlation decreases as the loss scale decreases, which indicates that the gradients of small errors trained on pseudo-labels are indeed noisier and less reliable. Therefore, it is reasonable to increase the confidence of the regression task by downweighting the gradients of small prediction errors.

### v2

@@ V-C1 Semi-supervised Setting
In addition to the existing semi-supervised landmark detection models [51, 22, 54, 56, 57], we also compare with semi-supervised pose estimation methods [33, 52] and generic self-training methods [16, 20, 29, 71]. For selection-based methods, we select their optimal thresholds based on the performance on 300W with 5% labeled data, which can be found in Fig. 8. Following [57], we also equip our model with HRNet [48] for better performance. Tab. I shows the results on 300W, WFLW, and AFLW with different labeled ratios. We can see that STLD obtains the best results on most of the entries. Specifically, STLD-HM-R18 and STLD-TF-R18 both outperform their ResNet-18 counterparts on almost all the entries, demonstrating the effectiveness of the proposed method. When a heavier backbone such as HRNet is equipped, the performance of STLD is further boosted. For example, STLD-HM-HRNet achieves 3.73, 3.40, and 3.32 of NME (%) on 300W for 1.6%, 5%, and 10% labeled data, respectively, outperforming all the previous methods. On WFLW and AFLW, STLD-HM-HRNet achieves the best results with 5% labeled data. In contrast, STLD-TF-HRNet performs better with higher labeled ratios, thanks to the representation capability of transformers. For example, it obtains 3.25 of NME (%) on 300W and 1.53 on AFLW with 20% labeled, both of which are slightly better than STLD-HM-HRNet. The advantage of STLD-TF-HRNet is even larger on WFLW, achieving 4.46 with 20% labeled, which is about 5.7% improvement over the best existing method HybridMatch (4.73). To make the evaluation comprehensive, we also include the results under AUC and FR10%, which are shown in Tab. II. Again, our STLD with HRNet obtains the SOTA performance on both metrics, validating the superiority of the proposed method.
| Method | Backbone | Unlabeled | MRE (mm) |
| SCN [76] | UNet | - | 0.66 |
| GU2Net [5] | UNet | - | 0.84 |
| HM† [69] | ResNet-50 | - | 0.68 |
| TF† [25] | ResNet-50 | - | 0.69 |
| Ours w/ HM | ResNet-50 | 12K | 0.64 |
| Ours w/ TF | ResNet-50 | 12K | 0.61 |
@@ V-D Model Analysis
| Method | 1.6% | 5% | 10% | 20% |
| No Pretrain | 4.26 | 3.64 | 3.48 | 3.34 |
| MAE [81] | 4.19(-1.6%) | 3.60(-1.1%) | 3.45(-0.9%) | 3.33(-0.3%) |
| PP (Ours) | 3.97(-6.8%) | 3.47(-4.7%) | 3.39(-2.6%) | 3.30(-1.2%) |
@@
How pseudo pretraining works. One may worry that the noisy samples learned during pseudo pretraining would be memorized by the model and thus leads to confirmation bias. To understand how it works, we perform an analysis of example forgetting by computing the difference of the pseudo-labels used for training and their predictions after the two-stage learning. Note that shrink regression is not used in the second stage here as we would like to focus on pseudo pretraining. As a comparison, we also do the analysis for the naive baseline. Fig. 10 shows the relevant results of both HM (top) and TF (bottom), with different labeled ratios, and the pseudo-labels are grouped by noise. First of all, we can see that the naive baseline forgets more about noisier samples, which is consistent to the findings from [82]. And as expected, the model with the two-stage training does not memorize the pseudo-labels as much as the naive baseline does because the pseudo-labels are not used in the second stage of the proposed method. But interestingly, the forgetting of noisier samples is enlarged when compared to the naive baseline (i.e., the pink area333The green dotted line of the pink area is obtained by moving the upper green solid line down and overlaps with the lower green solid line at the first black dot.), indicating that the two-stage design with pseudo pretraining is able to handle noisy pseudo-labels implicitly. Such a characteristic helps pseudo pretraining to effectively utilize the task-specific knowledge of pseudo-labels while avoiding the influence of noisy samples. To show the superiority of pseudo pretraining, we compare it with masked autoencoder (MAE) [81], a popular self-supervised pretraining method. For reference, we add a baseline named No Pretrain, which is directly fine-tuned on labeled data. Tab. V gives the results. We can see that pseudo-pretraining outperforms MAE consistently across all the labeled ratios, verifying the advantage of utilizing pseudo-labels over unlabeled pretraining as the former incorporates task-specific knowledge for better performance.
@@
How shrink loss works. While the shrink regression for heatmap-based models looks straightforward, the shrink loss applied to coordinate-based models seems not so obvious on adjusting task confidence. To verify it, we perform a correlation analysis of the gradients between GTs and pseudo-labels. We record the gradients444The seeds are fixed to make sure each pair of gradients from GTs and pseudo-labels is generated by the same batch of training data. of the last layer when trained with GTs and pseudo-labels, respectively, and group them by pseudo-label noise. For each group, we compute the Pearson correlation of the gradients between GTs and pseudo-labels. Fig. 11a and 11b show the results on 5% and 10% labeled settings of 300W, respectively. We can see that the correlation decreases as the loss scale decreases, which indicates that the gradients of small errors trained on pseudo-labels are indeed noisier and less reliable. Therefore, it is reasonable to increase the confidence of the regression task by downweighting the gradients of small prediction errors.
| Training | Inference | |||||
| Backbone | Self-training | STLD-HM | STLD-TF | Self-training | STLD-HM | STLD-TF |
| ResNet-18 | 0.75 hour | 1.00 hour | 6.00 hours | 3.25 ms | 3.34 ms | 2.61 ms |
| HRNet | 3.20 hours | 3.50 hours | 22.00 hours | 4.34 ms | 4.35 ms | 6.39 ms |
@@
Computational cost. Since the proposed method requires more training stages than a standard self-training method, the computational cost can be expensive. To investigate this, we compare the training time as well as inference time of our method with self-training [16]. We evaluate the total training time on 300W with 10% labeled data; for inference, we evaluate the inference time per image. Both training and inference were conducted with one RTX 3090 GPU. Table VI shows the computational cost of the two methods in different backbones for both training and inference. It can be seen that our STLD-HM model has a slightly longer training time than that of self-training for both ResNet-18 and HRNet, indicating that the training complexity of our method is acceptable as the two use the same base detector (i.e., heatmap regression). However, STLD-TF requires much longer training time (about 7 times), due to the slow convergence of transformers [27]. Such an issue might be mitigated by faster adaptations such as deformable detr [87], which spends 10 times less training epochs than the standard framework [27]. For inference speed, STLD-HM and self-training are again comparable. Surprisingly, STLD-TF with ResNet-18 is faster than the heatmap counterparts and we believe this is because coordinate regression does not require upsampling, which saves time. On the other hand, STLD-TF with HRNet is slower than the other two and we attribute it to the high-resolution feature map of HRNet, which significantly increases the computation of the self-attention in transformers.